40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 25 05:26:51 2011, kiwamu, Update, Green Locking, cavity scan cavity_scan.png
    Reply  Wed Jan 26 02:18:12 2011, Koji, Update, SUS, ETMX length to angle matrix 
    Reply  Wed Jan 26 10:11:47 2011, Aidan, Update, Green Locking, cavity scan 
Message ID: 4198     Entry time: Tue Jan 25 05:26:51 2011     Reply to this: 4204   4205
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: cavity scan 

cavity_scan.png

I scanned the X arm by changing an offset for the feedback to ETMX while the arm stayed locked by the green locking.

But the resultant plot is still far away from a beautiful one.

Changing the offset broke the lock frequently, so eventually I couldn't measure the stability of the IR-PDH signal at the resonance. 

 

 The plot above is a result of the scanning. You can see there is a clear resonance at the center of the plot.

However the lock frequently became unstable when I was changing the offset.

It looked like this unstability came from the end PDH lock. I guess there are two possible reasons:

  (1)  feedback range for the laser PZT is not wide enough. Right now the range is limited by a SR560, which has been used for a summing amplifier.

  (2)  Length to Alignment coupling. Pushing ETMX causes a misalignment.

The issue (1) can be easily solved by engaging the temperature feedback, which helps actuating the laser frequency a lot at DC.

The issue (2) will be also solved by well align the IR beam, the arm cavity and the green beam.

ELOG V3.1.3-